# Functional Clustering for Longitudinal Associations between Social Determinants of Health and Stroke Mortality in the US

### v2

@@ 3 Simulation
  • The parameter estimation accuracy is measured using the standardized mean square error (MSE) of the functional coefficients, which is defined as
    $\text{MSE}=\frac{\sum_{k=1}^{K}\sum_{j=1}^{p}\|\beta_{jk}-\hat{\beta}_{jk}\|_{2}^{2}}{\sum_{k=1}^{K}\sum_{j=1}^{p}\|\beta_{jk}\|_{2}^{2}}.$
@@
VS-RP 0.91 11.76 9.12 0.76 0.91 131.10 9.18 0.78 0.89 429.78 9.06 0.77 0.87 1176.21 9.18 0.79
LI-MIX 0.24 6.87 3.36 1.14 0.19 77.67 3.27 1.28 0.09 261.63 3.15 1.79 0.02 768.21 3.66 4.08
FS 0.99 11.99 0.25 0.09 0.98 131.69 0.33 0.09 1.00 430.90 0.05 0.06 0.96 1181.96 8.05 0.75
FS-Net 0.99 12.00 0.09 0.06 0.99 131.74 0.20 0.08 1.00 431.13 0.02 0.05 0.97 1181.95 5.90 0.50
FGS-Net 0.99 12.00 0.06 0.09 0.99 132.00 0.09 0.07 0.99 432.00 0.06 0.09 0.99 1182.00 0.66 0.12
We investigate performances of the above methods under various scenarios of $n$, $p$, and $\alpha$. Here, we set $n$ to $180$ or $300$, and take $p$ as $10,\ n/6,\ n/2,$ and $43n/$ 4, to consider the situations ranging from a small to a large number of covariates. Additionally, we set $\alpha$ to $0.4$ and $0.8$ to reflect the mild or strong dependence among the covariates. The averaged ARI, C, IC, and MSE are presented in Table 1. In the analysis below, we only focus on the results of $n=180$. Similar conclusions can be obtained from the result of $n=300$.
Overall, FGS-Net, FS-Net, and FS show superior performance under different scenarios of $n$, $p$, and $\alpha$, highlighting the advantages of implementing variable selection and clustering simultaneously under the REM framework. In contrast, the RP method only uses a roughness penalty and does not consider variable selections within the clustering. As a result, its performance quickly deteriorates for both clustering and parameter estimation as $p$ increases. Furthermore, among the two-step methods, we find that the VS-RP performs poorly compared to the first three REM-type methods. For instance, in all scenarios with $n=180$, the average ICs of VS-RP are mostly larger than 9, indicating that about half of the nonzero functional coefficients are incorrectly identified as zero. This leads to significant estimation errors in both clustering and parameter estimation by the VS-RP method, and suggests that selecting variables before clustering is ineffective. It’s worth noting that the results of LI-MIX are even worse, as this method further ignores the time-varying nature of $\bm{\beta}(\cdot)$ in the estimation procedure.
@@
In Table 1, we also observe that the ICs and MSEs of FS are significantly larger than those of FS-Net as $\alpha$ increases. This is expected since the dependencies between functional covariates may impede the performance of the FS procedure. This requires an additional ridge-type penalty in the FS-net to stabilize the estimation procedure. Moreover, as $p$ increases, the ICs and MSEs of FS-Net are further larger than those of FGS-net. For these cases, the high-dimensionality would undermine the statistical efficiency for both variable selection and clustering in the FS-Net. Therefore, it would be beneficial to impose cluster-invariant sparsity through FGS-net to borrow strengths across all clusters.
@@
Figure 3: The true curves of the functional coefficients $\beta_{jk}$ from each cluster $k$ and their estimations for $n=180$, $p=240$, and $\alpha=0.8$. The solid black line represents the true curve, while the solid red, green, and blue lines, along with their corresponding bands, represent the pointwise means and the 95% confidence bands over 100 replicated simulations.
@@
In the case of $n=180$, $p=240$, and $\alpha=0.8$, we further illustrate the estimation performance of the functional coefficients in Figure 3. We observe that the estimations of FS-Net exhibit smaller biases compared to FS, highlighting the significance of FS-Net in mitigating biases of functional coefficients. Furthermore, in comparison to those of FS-Net and FS, the estimated curves of FGS-Net show even smaller biases and narrower confidence bands, owing to the pursuit of cluster-invariant sparsity. Overall, FGS-Net is the most suitable choice among these six methods for clustering longitudinal associations and selecting important variables in high-dimensional functional covariates.
@@ 4 Real data
In this section, we apply our method to the SDOH and stroke mortality dataset for the clustering of their longitudinal associations. Given that the stroke mortality data are right-skewed and take positive values, we conduct a log transformation to stabilize its variance. Using our approach, we identify 2 clusters for the longitudinal associations and 18 relevant SDOH covariates for stroke mortality.
@@
The two clusters for the county-level longitudinal associations are presented in Figure 4. The proportions of two clusters, determined by the number of counties, stand at 68% and 32%, respectively. Notably, both clusters are prevalent across the majority of states in the US, encompassing both rural and urban areas (urban: 76% and 24%, and rural: 65% and 35%, for cluster 1 and cluster 2, respectively). It’s worth noting that the southeastern US contains a region called the Stroke Belt, known for its persistent high relative excess of stroke mortality. Despite counties in the Stroke Belt having similar stroke severity, this area is also mixed by the two clusters, with proportions of 70% and 30%, respectively. These results suggest that regions sharing similar geographic and stroke characteristics may have very different SDOH-stroke mortality associations, and we may need to consider separating two types of policies for the SDOH adjustments in stroke management based on our clustering results.
@@
Figure 4: The clustering result for the county-level longitudinal associations between the SDOH and stroke mortality in the US. The regions in grey color represent the area without available data.
@@
In addition, we illustrate the selected 18 covariates of SDOH in Table 4, ordered by their relative importance (Grömping, 2007) defined as
@@
$\text{RI}(\bm{\beta}_{j\cdot})=\|\bm{\beta}_{j\cdot}\|_{2}\left\{\frac{1}{n}\sum_{i=1}^{n}\big{\|}X_{ij}-\frac{1}{n}\sum_{i=1}^{n}X_{ij}\big{\|}^{2}\right\}^{1/2}.$
@@
We find that the influence of the SDOH on stroke mortality is mainly contributed by four aspects: social environment, built environment, health care system, and biology. Beyond the well-studied determinants from economic, cultural, and racial domains (Tsao et al., 2023), we find that stroke mortality is significantly associated with living and working environments, education level, and overuse of opioids. Typically, among the selected variables of SDOH, most of them are related to economic development. For example, in Table 4, MEDIAN_HOME_VALUE may reflect overall economic development and infrastructure building in the community. Additionally, a higher value of ELDERLY_RENTER may suggest a larger elderly population with lower income, facing issues such as housing instability. These economic-related factors may be potentially addressed through more equitable economic policies. For example, MEDIAN_HOME_VALUE can be adjusted by facilitating economic development. In addition, serving as a sign of the elderly living condition, ELDERLY_RENTER can be adjusted by improving elderly welfare.
@@
To further quantify the longitudinal associations between SDOH and stroke mortality, we investigate the functional coefficients $\beta_{jk}$ estimated from the dataset. Here, we only focus on the coefficients from the leading four important SDOH, which are MEDIAN_HOME_VALUE, NO_ENGLISH, ASIAN_LANG, and ELDERLY_RENTER. Their longitudinal associations in the two clusters are shown in Figure 5, along with their $95\%$ confidence bands obtained by bootstraps. We notice that the leading four SDOH possess time-varying associations with stroke mortality, which are distinguished by two clusters. Furthermore, these associations are notably different between the two clusters. For example, the fluctuation of the longitudinal associations in cluster 2 is more significant than those in cluster 1. Additionally, we observe that in cluster 2, all four associations exhibit a common inflection point in 2013 (the dotted vertical lines in Figure 5). These findings help us better understand the dynamics of the associations between SDOH and stroke mortality in US counties.
@@
In Figure 5, MEDIAN_HOME_VALUE shows a stable and weak association with stroke mortality for the counties in cluster 1, while those in cluster 2 exhibit a more intense negative effect with stroke mortality. The situation for ELDERLY_RENTER differs, with the association in cluster 2 showing significant fluctuation, while that in cluster 1 presents a time-increasing negative association with stroke mortality. Typically, economic factors have been found to affect stroke outcomes causally (Bann et al., 2023). As two economic-related factors, MEDIAN_HOME_VALUE and ELDERLY_RENTER have also been identified to possess significant associations with stroke mortality (Rodgers et al., 2019; Mawhorter et al., 2023). Our findings not only align with existing literature but also provide further insights into more tailored stroke mortality prevention strategies related to SDOH variables. For instance, the notable negative association between MEDIAN_HOME_VALUE and stroke mortality in cluster 2 highlights the importance of focusing on infrastructure development, maintenance, and enhancing overall economic equity in these regions. Noting the time-increasing negative association between ELDERLY_RENTER and stroke mortality in cluster 1, it may be beneficial to prioritize viable housing programs for lower-income elderly populations in these regions. Initiatives such as housing vouchers or subsidies may also be helpful, particularly in recent years when inflation is on the rise.
@@
In addition to the above two covariates, NO_ENGLISH and ASIAN_LANG are two other SDOH variables measuring the sociocultural environment of the county. NO_ENGLISH, an indicator of the percentage of the population that does not speak English at all, is found to be associated with stroke mortality in both clusters. The direction of the associations shifts from positive to negative. A similar decreasing trend is observed in the association between ASIAN_LANG (a measure of the percentage of the population that speaks Asian and Pacific Island languages) and stroke mortality in cluster 2. Meanwhile, the association between ASIAN_LANG and stroke mortality in cluster 1 remains stable and weakly positive over time.
@@
Our result shows that the association between the density of immigrants or Asian and Pacific Islanders and stroke mortality may not be uniform across regions and time periods. It is worth considering that immigrants or Asian and Pacific Islanders who experience a stroke may encounter challenges in accessing timely stroke care if they do not speak English. This language barrier may potentially contribute to an increased risk of stroke-related death in specific regions and during certain time periods (Shah et al., 2015). However, as suggested by our results, their risks with stroke-related death may have diminished from 2010 to 2018. This implies a gradual improvement of stroke care for immigrants or Asian and Pacific Islanders in the US, particularly for the counties in cluster 2.
@@ 5 Dicussion
In this article, we introduce a novel clustering method for regional divisions of US counties based on their longitudinal associations between SDOH and stroke mortality. The challenges for this task arise from the latent and cluster-specific nature of the associations, which are compounded by their functional and high-dimensional characteristics simultaneously. To tackle these complex structures, we propose an REM algorithm that utilizes a finite mixture of functional linear concurrent models. Our method explores the clustering-invariant sparsity via a FGSCAD-Net penalty within the REM algorithm, allowing for efficient variable selection in the functional covariates to identify the most significant associations among all clusters. The effectiveness of our method has been demonstrated via extensive simulations. In the end, we apply the proposed method to our SDOH – stroke mortality longitudinal data, facilitating the regional divisions of US counties. This enables the identification of regions for informing SDOH–targeted prevention of stroke mortality.
In the analysis of the SDOH and stroke mortality dataset, our clustering map represents a novel result for region-division in stroke management, taking into account the similarity among the longitudinal associations between SDOH and stroke mortality. These findings suggest that heterogeneity in the associations occurs even within areas sharing similar geographical conditions, stroke mortality, or economic status. This implies that the region divisions solely based on the above factors may not be effective for SDOH-based stroke death control. Moreover, we uncover various patterns of longitudinal associations among the two identified clusters in US counties. The dynamics within these associations, including scale, trends, and inflection points, not only provide heuristic information for understanding complex SDOH – stroke mortality associations but are also useful for establishing timely SDOH adjustments in stroke death control. Finally, the proposed clustering method can also be generalized to other high-dimensional longitudinal datasets, serving for clustering and variable selections in their longitudinal associations.
SUPPLEMENTARY MATERIAL
“SuppMaterial.pdf”:
This file describes the details related to the proposed approach method, including the dataset, the selection scheme of tuning parameters, and other supporting results of the simulation study.

### v6

@@ 3.1 Data Generation
  • The parameter estimation accuracy is measured using the standardized mean square error (MSE) of the functional coefficients: $\text{MSE}=\frac{\sum_{k=1}^{K}\sum_{j=1}^{p}\lVert\beta_{jk}-\hat{\beta}_{jk}\rVert_{2}^{2}}{\sum_{k=1}^{K}\sum_{j=1}^{p}\lVert\beta_{jk}\rVert_{2}^{2}}.$
3.2 Result
We investigate the performances of different methods under various scenarios of $n$, $p$, and $\alpha$. Here we set $n$ to $180$ or $300$ and take $p$ as $210,n/6,n/$ 2, and $43n/$ 4, to consider the situations ranging from a small to a large number of covariates. Additionally, we set $\alpha$ to $0.4$ and $0.8$ to reflect the mild or strong dependence among the covariates. The averaged ARI, C, IC, and MSE are presented in Table 1. In the analysis below, we only focus on the results of $n=180$. Similar conclusions can be obtained for the result of $n=300$.
Overall, FGS-Net, FS-Net, and FS demonstrate superior performance across various scenarios with different values of $n$, $p$, and $\alpha$, underscoring the benefits of simultaneously implementing variable selection and clustering under the REM framework. In contrast, the RP method, which only applies a roughness penalty and omits variable selection within the clustering process, quickly deteriorates in performance for both clustering and parameter estimation as $p$ increases. Among the two-step methods, we find that VS-RP performs poorly compared to the first three REM-type methods. For instance, in all scenarios with $n=180$, the average ICs of VS-RP are mostly larger than 9, indicating that about half of the nonzero functional coefficients are incorrectly identified as zero. These results lead to significant estimation errors in both clustering and parameter estimation by the VS-RP method, suggesting that selecting variables prior clustering is ineffective. Notably, the performance of LI-MIX is even poorer, as this method additionally ignore the time-varying nature of $\bm{\beta}(\cdot)$ during the estimation process.
@@ 3.2 Result
VS-RP 0.92 11.73 8.04 0.7 0.91 130.92 7.89 0.7 0.9 429.57 7.74 0.68 0.85 1175.85 8.13 0.69
LI-MIX 0.24 6.42 2.88 1.16 0.2 80.13 2.91 1.26 0.08 263.16 3.3 1.82 0.02 747.66 3.57 4.57
FS 1.00 12.00 0.00 0.04 1.00 131.65 0.00 0.08 0.99 430.79 0.09 0.04 0.97 1181.97 8.02 0.72
FS-Net 1.00 11.99 0.00 0.04 1.00 131.74 0.00 0.04 0.99 431.05 0.07 0.04 0.98 1181.96 6.14 0.49
FGS-Net 0.99 12.00 0.00 0.04 0.99 132.00 0.06 0.04 1.00 432.00 0.00 0.04 0.99 1182.00 0.30 0.09
@@
In Table 1 we observe that the ICs and MSEs of FS are significantly larger than those of FS-Net as $\alpha$ increases. This is expected since the dependencies between functional covariates may impede the performance of the FS procedure. Incorporating an additional ridge-type penalty in FS-Net could help stabilize the estimation process for both variable selection and parameter estimation. Moreover, as $p$ increases, the ICs and MSEs of FS-Net become further larger than those of FGS-Net. In these scenarios, the high dimensionality may undermine the statistical efficiency for both variable selection and clustering in FS-Net. Therefore, it would be beneficial to impose cluster-invariant sparsity through FGS-net to borrow strength across all clusters during estimation.
@@
Figure 3: The true curves of the functional coefficients $\beta_{jk}$ from each cluster $k$ and their estimations from FS, FS-Net, and FGS-Net, for $n=180$, $p=240$, and $\alpha=0.8$. The solid black lines represents the true curve, while the dashed lines, along with their corresponding bands, represent the pointwise means and the 95% confidence bands over 100 replicated simulations.
@@
In the case of $n=180$, $p=240$, and $\alpha=0.8$, we further illustrate the estimation performance of the functional coefficients in Figure 3. We observe that the estimations of FS-Net exhibit smaller biases, compared to FS, highlighting the significance of FS-Net in mitigating biases of functional coefficients. In comparison to the results of FS-Net and FS, the estimated curves of FGS-Net show even smaller biases and narrower confidence bands, owing to the pursuit of cluster-invariant sparsity.
In Section C.4 of the Supplementary Material, we assess the computational cost of FGS-Net in comparison to the two-step methods VS-RP and LI-MIX. Our findings indicate that, while the two-step methods are computationally faster than those utilizing REM algorithms, their performance is generally inferior, as shown in Table 3, due to the uncorrelatedness of the variable selection and clustering steps. Additionally, we observe that the computational cost of FGS-Net does not increase too rapidly as $n$ and $p$ increase. This suggests tolerable computational costs for simultaneously clustering and variable selection in longitudinal associations for large datasets.
@@ 4 Real data
In this section we apply our methods to the SDOH and stroke mortality dataset for clustering their longitudinal associations. Given that the stroke mortality data are right-skewed and take positive values, we conduct a log transformation to stabilize its variance. Besides, we normalize the data $X_{ij}(t_{S_{i}})$ by $({\sum_{i}S_{i}})^{-1}\sum_{i=1}^{n}\sum_{s=1}^{S_{i}}X_{ij}(t_{S_{i}})^{2}$ to ensure that the SDOH covariates are on the same scale.
Performance Comparison
We apply the proposed FS-Net and FGS-Net to the dataset and compare these methods with FMFLCM(1) (Wang et al. (2008); Goldsmith and Schwartz (2017); Ghosal et al. (2020)) and the LI-MIX method (Peel and MacLahlan (2000); Khalili and Chen (2007); Khalili and Lin (2013)) in the simulation section. Notably, FMFLCM(1) is a special case of the FMFLCM in (1) with $K=1$, indicating that it does not cluster longitudinal associations during variable selection. Furthermore, LI-MIX does not account for longitudinal changes in the SDOH-stroke mortality associations during clustering, and FS-Net focuses only on cluster-variant sparsity in model estimation. Detailed comparisons of these methods are presented in Section C.3 of the Supplementary Material. Overall, FMFLCM(1) and LI-MIX tend to select a large number of covariates but fail to adequately fit the data. In contrast, FS-Net and FGS-Net provide a more accurate fit, highlighting the importance of accommodating longitudinal changes and heterogeneity when analyzing associations between SDOH and stroke mortality.
It is worth noting that the model in FS-Net has a higher complexity, compared to that of FGS-Net, with the latter pursuing cluster-invariant sparsity rather than cluster-variant sparsity. In our analysis we find that both methods achieved the fitting residuals of 0.15, and their performances are quite similar in terms of clustering and variable selection. These findings suggest that the increased complexity of FS-Net does not result in significant improvements in model fitting. For a more interpretable result, we focus only on the data analysis of FGS-Net in the remaining and refer to Section C.3 of the Supplementary Material for the results of FS-Net.
Data Analysis
@@
Using FGS-Net, we identify two clusters for the longitudinal associations and 18 relevant SDOH covariates for stroke mortality. The two clusters for the county-level longitudinal associations are presented in Figure 4. The proportions of two clusters, determined by the number of counties, stand at 68% and 32%, respectively. Notably, both clusters are prevalent across the majority of states in the U.S., encompassing both rural and urban areas (urban: 76% and 24%, and rural: 65% and 35%, for cluster 1 and cluster 2, respectively). It is worth noting that the southeastern U.S. contains a region called the Stroke Belt (Heyden et al. (1978); Lanska and Kuller (1995); Karp et al. (2016); Howard and Howard (2020)), known for its persistent high relative excess of stroke mortality. Despite counties in the Stroke Belt having similar stroke severity, this area is also mixed by the two clusters, with proportions of 70% and 30%, respectively. These results suggest that regions sharing similar geographic and stroke characteristics may have very different SDOH-stroke mortality associations, and we may need to consider separating two types of policies for the SDOH adjustments in stroke management based on our clustering results.
@@
Figure 4: The clustering result for the county-level longitudinal associations between the SDOH and stroke mortality in the US.
@@
In addition, we illustrate the selected 18 covariates of SDOH in Table 4, ordered by their relative importance (Grömping (2007)), defined as
@@
$\text{RI}(\bm{\beta}_{j\cdot})=\lVert\bm{\beta}_{j\cdot}\rVert_{2}\Biggl{\{}\frac{1}{n}\sum_{i=1}^{n}\Biggl{\lVert}X_{ij}-\frac{1}{n}\sum_{i=1}^{n}X_{ij}\Biggr{\rVert}^{2}\Biggr{\}}^{1/2}.$
@@
We find that the influence of the SDOH on stroke mortality is mainly contributed by four aspects: social environment, built environment, health care system, and biology. Beyond the well-studied determinants from economic, cultural, and racial domains (Tsao et al. (2023)), we find that stroke mortality is significantly associated with living and working environments, education level, and overuse of opioids. Typically, among the selected variables of SDOH, most of them are related to economic development. For example, in Table 4, MEDIAN_HOME_VALUE may reflect overall economic development and infrastructure building in the community. Additionally, a higher value of ELDERLY_RENTER may suggest a larger elderly population with lower income, facing issues such as housing instability. These economic-related factors may be potentially adjusted through more equitable economic policies. For example, MEDIAN_HOME_VALUE can be adjusted by facilitating economic development. In addition, serving as a sign of the elderly living condition, ELDERLY_RENTER can be adjusted by improving elderly welfare.
@@
To further quantify the longitudinal associations between SDOH and stroke mortality, we investigate the functional coefficients $\beta_{jk}$ estimated from the dataset. Here we only focus on the coefficients from the leading four important SDOH, which are MEDIAN_HOME_VALUE, NO_ENGLISH, ASIAN_LANG, and ELDERLY_RENTER. The longitudinal associations in the two clusters, along with the 95% confidence bands, are shown in Figure 5. These 95% confidence bands are constructed using a wild-bootstrap procedure similar to those in Zhu et al. (2012). The detailed bootstrap procedure and its effectiveness are demonstrated in Section C.2 of the Supplementary Material.
In Figure 5 we observe that the leading four SDOH possess time-varying associations with stroke mortality. Furthermore, these associations are notably different between the two clusters. For example, the fluctuation of the longitudinal associations in cluster 2 is more significant than those in cluster 1. Additionally, we observe that, in cluster 2, all four associations exhibit a common inflection point in 2013 (the dotted vertical lines in Figure 5). These findings help us better understand the dynamics of the associations between SDOH and stroke mortality in the U.S..
In Figure 5, MEDIAN_HOME_VALUE shows a stable and weak association with stroke mortality for the counties in cluster 1, while those in cluster 2 exhibit a more intense negative effect with stroke mortality. The situation for ELDERLY_RENTER differs with the association in cluster 2 showing significant fluctuation, while that in cluster 1 presents a time-increasing negative association with stroke mortality.
@@
As economic-related factors, MEDIAN_HOME_VALUE and ELDERLY_RENTER have been identified to significantly associate with stroke mortality (Rodgers et al. (2019); Mawhorter et al. (2023)). Our findings not only align with existing literature but also provide further insights into more tailored stroke mortality prevention strategies related to SDOH variables. For instance, the notable negative association between MEDIAN_HOME_VALUE and stroke mortality in cluster 2 highlights the importance of focusing on infrastructure development, maintenance, and enhancing overall economic equity in these regions. Noting the time-increasing negative association between ELDERLY_RENTER and stroke mortality in cluster 1, it may be beneficial to prioritize viable housing programs for lower-income elderly populations in these regions.
@@
In addition, NO_ENGLISH and ASIAN_LANG are SDOH variables that measure the sociocultural environment of the county. NO_ENGLISH, which indicates the percentage of the population that does not speak English, is associated with stroke mortality in both clusters, with the direction of the associations shifting from positive to negative. A similar decreasing trend is observed in the association between ASIAN_LANG (measuring the percentage of the population that speaks Asian and Pacific Island languages) and stroke mortality in cluster 2. Conversely, the association between ASIAN_LANG and stroke mortality in cluster 1 remains stable, exhibiting a weakly positive trend over time.
@@
The above results suggest that the association between the density of immigrants or Asian and Pacific Islanders and stroke mortality may not be uniform across regions and time periods. It is worth considering that immigrants or Asian and Pacific Islanders who experience a stroke may encounter challenges in accessing timely stroke care if they do not speak English. This language barrier may potentially contribute to an increased risk of stroke-related death in specific regions and during certain time periods (Shah et al. (2015)). However, our findings indicate that their risks with stroke-related death may have diminished from 2010 to 2018. This reduction may be caused by gradual improvements in stroke care for immigrants or Asian and Pacific Islanders in the U.S., particularly in the counties of cluster 2.
@@ 5 Dicussion
In this article, we introduce a novel clustering method for regional divisions of U.S. counties based on their longitudinal associations between SDOH and stroke mortality. The challenges for this task arise from the latent and cluster-specific nature of the associations, which are compounded by their functional and high-dimensional characteristics simultaneously. To tackle these complex structures, we propose an REM algorithm that utilizes a finite mixture of functional linear concurrent models. Our method explores the clustering-invariant sparsity via a FGS-Net penalty within the REM algorithm, allowing for efficient variable selection in functional covariates to identify the most significant associations among all clusters. The effectiveness of our method has been demonstrated via extensive simulations. In the end we apply the proposed method to the SDOH–stroke mortality longitudinal data, facilitating the regional divisions of U.S. counties. This enables the identification of regions for informing SDOH–targeted prevention of stroke mortality.
In the analysis of the SDOH and stroke mortality dataset, our clustering map in Figure 4 provides a novel result for region-division in stroke management, taking into account the similarity among longitudinal associations between SDOH and stroke mortality. These findings indicate that heterogeneity in these associations occurs even within areas that share similar geographical conditions, stroke mortality rates, or economic statuses. Therefore, region divisions based solely on these factors may not be effective for SDOH-based stroke death control. Moreover, we uncover various patterns of longitudinal associations among the two identified clusters of U.S. counties. The dynamics within these associations, including scale, trends, and inflection points, not only provide heuristic information for understanding complex SDOH-stroke mortality associations but also are useful for establishing timely SDOH adjustments in stroke death control.
Our method can be generalized to other clustering tasks, such as time-varying clustering (Sartório and Fonseca (2020)) and spatial smooth clustering (Li and Sang (2019); Liang et al. (2021)). In these applications, we can relax the assumption of fixed cluster membership to accommodate the desired clustering structures. For instance, to implement time-varying clustering, we can allow cluster memberships to change over time within the finite mixture of functional linear concurrent models; for spatial smooth clustering, spatial dependence structures can be integrated into the model of cluster memberships (Jiang and Serban (2012); Liang et al. (2021)). All these extensions can be achieved by modifying the likelihood function in the REM algorithm, tailored to specific research problem of interest. We leave these as future directions for investigation.